OpenStudy (anonymous):

Does anyone know of a simpler way to solve the system the substitution method? The following is a problem I have on hand; x-3y=-15 4y-6x=62

OpenStudy (unklerhaukus):

multiply the first equation by 6 add this to the second equation, the x's will cancel out

OpenStudy (anonymous):

So confirming it will look something like this? (6x)-3y=-15

OpenStudy (agent0smith):

You have to multiply the whole first equation by 6. first rearrange the order like so (change the order of the second equation): x - 3y = -15 -6x + 4y = 62 Now multiply everything in the first equation by 6, on both sides of the equals sign: 6 ( x - 3y = -15 ) -6x + 4y = 62

OpenStudy (unklerhaukus):

x-3y =-15 (1st Eqn) 6(x-3y) = 6(-15) (1st Eqn times 6) =

OpenStudy (anonymous):

That makes sense now AgentO... Now will this give me the answer for x or y?

OpenStudy (agent0smith):

It'll give you y, since it'll eliminate x. 6 ( x - 3y = -15 ) -6x + 4y = 62 becomes 6x - 18y = -90 -6x + 4y = 62 Now add the two equations together.

OpenStudy (anonymous):

now when you say add the two equations together, how does that look?

OpenStudy (unklerhaukus):

[6x - 18y] = [-90] {-6x + 4y} = {62} [6x - 18y ]+{-6x + 4y } = [-90] +{62} 6x - 18y -6x + 4y = -90 + 62

OpenStudy (anonymous):

so is the final answer is (6,2) ?

OpenStudy (agent0smith):

Adding them vertically is easier imo, since they're already lined up. 6x - 18y = -90 -6x + 4y = 62 -------------- 0 - 14y = -28

OpenStudy (anonymous):

That is easier, thx for the tip...

OpenStudy (unklerhaukus):

so you've got y=2 right,

OpenStudy (unklerhaukus):

sub this into the first equation to get x

OpenStudy (anonymous):

So instead of it being x-3y=-15 it would be x-3(2)=-15 to get x?

OpenStudy (unklerhaukus):

correct ..

OpenStudy (anonymous):

Got it... (-9,2) Thank you both for ur assistance. this was a big help.

OpenStudy (unklerhaukus):

that's it \[\boxed{\large\color{red}\checkmark}%unk\]

OpenStudy (anonymous):

Awesome.

OpenStudy (unklerhaukus):

and you can always check your solution by subbing the solution into the original equations x-3y=-15 4y-6x=62 (x,y)=(-9,2) (-9)-3(2)=-15 4(2)-6(-9)=62 both these should be true

OpenStudy (agent0smith):

^ learning to check your solutions is a good idea. You never need anyone to check your work (on systems of equations) if you check your own answers.

OpenStudy (anonymous):

that's a cool way to check ur work , thx for the tip.